OpenStudy (anonymous):

i need help with a problem: √11x =3√x+2

OpenStudy (lalaly):

square both sides 11x=9x+2 2x=3 x=1

OpenStudy (lalaly):

2x=2

OpenStudy (lalaly):

sorry

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{11x}=3\sqrt{x+2}\] \[\large (\sqrt{11x})^2=(3\sqrt{x+2})^2\] \[\large 11x=9(x+2)\] \[\large 11x=9x+18\] \[\large 11x-9x=18\] \[\large 2x=18\] \[\large x=\frac{18}{2}\] \[\large x=9\]